OpenStudy (anonymous):

2v^2+7v+5 Factor by grouping I am stuck on this one too! Help

OpenStudy (anonymous):

Is it (v+1) (2v+5)?

OpenStudy (anonymous):

Expand that out and see if you get 2v^2+7v+5

OpenStudy (anonymous):

If you get that, then it's correct.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

worked! thanks

OpenStudy (anonymous):

how about 3x-5=7x+11

OpenStudy (anonymous):

trying to find solution set

OpenStudy (anonymous):

Uh, this is much simpler than what you did before. Just collect the like terms and you can find what x is....

OpenStudy (anonymous):

What's so hard about that one?

OpenStudy (anonymous):

but its not -4

OpenStudy (anonymous):

its sets

OpenStudy (anonymous):

like a quadratic I think

OpenStudy (anonymous):

3x-5=7x+11 How is that a quadratic?

OpenStudy (anonymous):

A QUADRATIC MUST HAVE AN x^2!!!!

OpenStudy (anonymous):

well its asking for sets which mean two different answers right??

OpenStudy (anonymous):

sory about the quadratic

OpenStudy (anonymous):

Can't you write the whole question out word for word please? Because it's confusing to just look at the equation and instantly know what to do....

OpenStudy (anonymous):

let me try it just as solving but it asks for the solution set

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve the equation 3x-5=7x+11 the solution set is { }

OpenStudy (anonymous):

What's the answer?

OpenStudy (anonymous):

on your answer sheet.

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

doesnt have one thats the problem

OpenStudy (anonymous):

Then what's so bad about that when you know how to get -4 and the answers say -4?

OpenStudy (anonymous):

-4 is the answer. What's so problematic about this?

OpenStudy (anonymous):

not sure but thinking it is different answer. Lets see here goes.